# Integrating Mechanistic Modeling and Machine Learning to Study CD4+/CD8+ CAR-T Cell Dynamics with Tumor Antigen Regulation

### v1

@@ 3.2 Model Validation
To further assess the impact of CD4:CD8 treatment ratios on therapeutic efficacy, we simulated additional treatment conditions corresponding to various fractions of CD4 cells, calculating the percentage of virtual patients who improved from non-responder (NR) to complete responder (CR) under each ratio (5). Our simulations demonstrated a higher rates of treatment response with increasing CD4:CD8 ratios, up to a certain extent (the presence of CD8 T cells increased treatment outcomes); these results are in agreement with in vivo data showing the same CAR-T expansion patterns with varying CD4:CD8 ratios (Lee et al. 2018).
Figure 5: Fig. 5 Percentage of virtual patients who respond to treatment (switch from NR to CR when treatment is applied) across different CD4:CD8 treatment ratios. Higher CD4 fractions in the infused CAR-T therapy led to greater rates of tumor clearance, consistent with prior findings that increased CD4:CD8 ratios enhance CAR-T proliferation and therapeutic efficacy (Lee et al. 2018)
The CD4:CD8 ratio in CAR-T therapy without a defined formulation can vary widely depending on the patient’s original lymphocyte composition. Galli et al analyzed the CD4:CD8 ratios in such patients, and found that on average, complete responders had a CD4:CD8 ratio that was closer to 1:1 (smaller) than non-responders (Galli et al. 2023). To test our model against this finding, the virtual patients were treated with a random CD4:CD8 ratio (fraction of CD4 cells was randomly chosen from 0 to 1), and then categorized as CR or NR. It was found that the average fraction of cells that were CD4 in complete responders was 0.488 (standard deviation 0.268), and in nonresponders it was 0.558 (standard deviation 0.301). This is in agreement with the data from Galli et al. (2023); on average, patients who respond to treatment have a higher CD4:CD8 ratio (closer to 1:1).
@@ 3.3 Parameter Selection for Practical Modeling
In theory, knowing the values of a patient’s parameters would allow prediction of treatment outcomes for any CD4:CD8 ratio using the model. In practice, most of these parameters, although biologically meaningful, are impractical or impossible to measure before treatment. However, some of the most sensitive parameters ($N$, $k_{b2}$, $k_{e}^{CD8}$, $k_{b1}$, $B50_{CD8}$) can be approximated as follows.
@@ 3.4 Neural Networks for Treatment Response Prediction
If the five aforementioned parameters can be measured, they can be used as inputs to the model, while all other parameters are set to typical patient values, allowing simulation of treatment outcomes. To evaluate the robustness of our tumor-immune mathematical model to parameter uncertainty, we generated 5,000 virtual patients by sampling patient-specific values for the five aforementioned parameters from known physiological ranges (Kirouac et al. 2023). For each patient, we then created a corresponding noisy parameter set by drawing values from a normal distribution centered at the true parameter values, with standard deviations proportional to the width of each parameter’s range. Each patient was simulated under multiple treatment scenarios (no treatment, and CAR-T cell therapies with varying CD4:CD8 ratios) using both the true and noisy parameter sets. We classified treatment outcomes (complete response vs. non-response) for both parameter sets and assessed the consistency between them to quantify the model’s robustness.
Treatment TPR (%) TNR (%) FPR (%) FNR (%) Bal. Acc (%)
NoTx 30.6 31.9 26.4 11.1 31.3
CD4frac 0.0 35.4 26.8 27.2 10.6 31.1
CD4frac 0.5 35.6 26.7 27.2 10.4 31.2
CD4frac 0.8 34.1 28.0 27.3 10.7 31.1
CD4frac 1.0 32.1 30.2 26.8 10.9 31.2
@@
Table 3: Robustness of the tumor–immune ODE model to parameter noise ($n=5000$). Metrics shown are true positive rate (TPR, sensitivity), true negative rate (TNR, specificity), type I error rate (FPR), type II error rate (FNR), and balanced accuracy (mean of TPR and TNR). Classification is CR vs. NR under baseline versus noisy parameter sets across CD4:CD8 treatment conditions.
@@
Metrics such as true positive rate (sensitivity), true negative rate (specificity), type I error rate (false positives), and type II error rate (false negatives) were computed across all treatment conditions to summarize performance under noisy input data (Table 3). The average balanced accuracy (mean of sensitivity and specificity) across treatments was about 31.1%; this shows that the accuracy of the model was extremely compromised when incorporating noise into the parameter values. This noise, which reflects real-world variability in biological measurements, makes it difficult to reliably predict treatment outcomes. To address this challenge, we implemented a feedforward neural network to train the model on noisy data. The goal was to leverage the network’s ability to learn patterns from noisy inputs and improve the model’s predictive accuracy, thus mitigating the impact of measurement noise on the treatment outcome predictions.
@@
We trained a multi-label feed-forward neural network to predict, for each virtual patient, whether the mechanistic model yields a complete response (CR) or non-response (NR) under five treatment conditions simultaneously (one output per treatment). Inputs were the patient’s noisy values for five model parameters; targets were the corresponding noiseless simulation outcomes (CR vs. NR). Each input variable was standardized by subtracting the training-set mean and dividing by the training-set standard deviation to place features on a comparable scale.
@@
Data were split into a held-out test set (20%) and, from the remaining 80%, a validation set (20% of the remainder; 16% overall) and a training set. Training ran for up to 200 epochs with early stopping based on validation balanced accuracy at a fixed 0.5 threshold, restoring the best checkpoint. After training, decision thresholds were calibrated per treatment by sweeping a grid $t\in[0.01,0.99]$ on the validation set and selecting the threshold that maximized validation balanced accuracy. These validation-optimized thresholds were then applied once to the test set, and we report per-treatment sensitivity, specificity, type I/II error, and balanced accuracy (Table 4).
Treatment TPR (%) TNR (%) FPR (%) FNR (%) Bal. Acc (%)
NoTx 66.0 66.5 33.5 34.0 66.3
CD4frac 0.0 73.5 60.8 39.2 26.5 67.2
CD4frac 0.5 72.5 60.7 39.3 27.5 66.6
CD4frac 0.8 74.2 57.4 42.6 25.8 65.8
CD4frac 1.0 77.3 54.8 45.2 22.7 66.0
Table 4: Performance of the neural network in classifying treatment response (CR vs. NR) on the test set (20% of the 5,000 virtual patients) across CD4:CD8 treatment conditions. Metrics shown are true positive rate (TPR, sensitivity), true negative rate (TNR, specificity), type I error rate (FPR), type II error rate (FNR), and balanced accuracy (mean of sensitivity and specificity).
@@
As epochs increase, the training loss steadily decreases, while the validation loss drops at first and then plateaus; stopping training near the point where the validation loss plateaus (Fig. 6) helps prevent overfitting; i.e., it keeps the model from learning dataset-specific noise and preserves generalization to unseen data. In Fig. 7, the validation balanced accuracy rises quickly over the early epochs and then levels off, indicating that the network is learning useful signal and then converging. The plateau (with only small fluctuations) suggests diminishing returns from additional training rather than continued improvement, consistent with a stable model that generalizes to held-out data. Because balanced accuracy averages sensitivity and specificity, the upward trend reflects gains in both detecting responders and correctly rejecting non-responders despite class imbalance.
The mean balanced accuracy across treatments was 66.4%, compared with 31.1% for direct input into the model (Table 3). The neural network demonstrates a moderate ability to predict treatment outcomes, if not a significant improvement from directly use of the model.
@@
We used SHAP analysis to interpret feature importance across treatments (Figure 8(a)). Consistent with prior sensitivity analyses, variability in $N$ was the most influential factor overall (Figure 8(b)). The relative importance of the remaining parameters ($k_{b2}$, $k_{e}^{CD8}$, $k_{b1}$, $B50_{CD8}$) varied slightly across local, PRCC, and Sobol analyses, which is reflected in the SHAP ranking not exactly matching the Sobol order (Figure 8(c)). This indicates that while $N$ largely drives system behavior, interactions among the other parameters likely contribute nonlinearly to treatment outcomes. These results illustrate the value of combining mechanistic modeling with machine learning to better understand complex parameter effects under uncertainty.
@@ 4 Discussion
In this study, we extended the CAR-T cell therapy model proposed by Kirouac et al., which models antigen-regulated transitions between memory, effector, and exhausted T cell states, to explicitly model CD4+ and CD8+ T cell subtypes and their distinct roles in tumor-immune dynamics. As far as we know, our model is the first to differentiate between CD4+ and CD8+ CAR-T cells, allowing investigation of the synergistic interactions between these subtypes and their impact on treatment outcomes.
@@
We used our extended model to investigate the effects of varying CD4:CD8 ratios on CAR-T cell efficacy and tumor control. Specifically, we simulated treatment scenarios with different CD4:CD8 combinations, comparing outcomes in terms of tumor burden and response rates. Our results demonstrate that a balanced 1:1 ratio of CD4+ to CD8+ CAR-T cells leads to more effective tumor elimination on average (compared with a nonspecific CD4:CD8 ratio), consistent with clinical data showing enhanced CAR-T cell expansion and improved therapeutic outcomes at this ratio. As such, an important next step would be identification of the patient factors that determine whether a 1:1 CD4:CD8 treatment ratio or an unspecified ratio yields better therapeutic outcomes. Machine learning can be employed to analyze virtual patient simulation results to identify these factors. Identification of patient subgroups that are most likely to benefit from a specific CD4:CD8 ratio can enhance patient stratification and allow more personalized treatment.
@@
Additionally, our model allowed for exploration of how patient-specific variations in immune and tumor characteristics influence response to treatment. Although noise in parameter measurements may hinder direct predictions of treatment outcomes using the mathematical model, machine learning provides a potential solution for identifying which patients are likely to respond. We trained a neural network on noisy parameter sets and corresponding simulation outcomes to predict virtual patient responses, achieving a balanced accuracy of 66.4% (a significant improvement from 31.1% for direct predictions from the mathematical model). Additionally, the agreement between SHAP and mechanistic sensitivity analyses strengthens confidence in the model’s predictions and illustrates the complementarity of mechanistic and machine learning approaches.

### v2

@@ 3.2.1 Consistency of model outcomes with reported effects of 1:1 CAR-T composition
3.2.2 Response and expansion behavior across CD4:CD8 ratios
[…]
Figure 5: Fig. 5 Percentage of baseline non-responders rescued to complete response as a function of CD4 fraction in the CAR-T product. Results are shown for five independent trials of 5,000 virtual patients each, with newly generated virtual patient cohorts in every trial. Bars represent the mean rescue percentage, and error bars denote the standard deviation across trials
Figure 6: Fig. 6 CAR-T fold expansion as a function of CD4 fraction during the early and late phases of the initial expansion (“burst”) period. Shown is the mean fold expansion at $t=7$ (top) and $t=21$ (bottom) across simulated virtual patients for varying CD4 fractions in the infused CAR-T product. Fold expansion was quantified as the total CAR-T cell population evaluated at the indicated time point, normalized by the cumulative effective treatment exposure, computed as the area under the curve (AUC) of the DoseX state from infusion to the evaluation time. Error bars denote the standard deviation across patients. Results are averaged over five independent trials, each consisting of 5,000 independently generated virtual patients
Second, because the experimental study by Lee et al. primarily reported CAR-T expansion rather than response classifications, we directly examined CAR-T expansion dynamics as a function of CD4 fraction (Figure 6) over time. We computed the mean fold expansion of the total CAR-T population during the early ($t=7$) and later ($t=21$) phases of the post-infusion expansion burst. Early expansion was modestly higher for CD8-enriched products, consistent with the dominant cytotoxic role of CD8+ effector cells, whereas at later times increased CD4 fractions supported slightly greater overall expansion. Notably, the observed trends depend on the time at which expansion is evaluated, and large error bars indicate pronounced inter-patient variability; accordingly, these results are interpreted as broadly qualitative rather than definitive.
3.2.3 Response stratification under random CD4:CD8 CAR-T compositions
The CD4:CD8 ratio in CAR-T therapy without a defined formulation can vary widely depending on the patient’s original lymphocyte composition. Galli et al. analyzed the CD4:CD8 ratios in such patients, and found that on average, complete responders had a CD4:CD8 ratio that was closer to 1:1 (smaller) than non-responders (Galli et al. 2023). To test our model against this finding, the virtual patients were treated with a random CD4:CD8 ratio (fraction of CD4 cells was randomly chosen from 0 to 1), and then categorized as CR or NR. It was found that the average fraction of cells that were CD4 in complete responders was 0.488 (standard deviation 0.268), and in non-responders it was 0.558 (standard deviation 0.301). This is in agreement with the data from Galli et al. (2023); on average, patients who respond to treatment have a higher CD4:CD8 ratio (closer to 1:1).
@@ 3.3 Parameter Selection for Practical Modeling
In theory, knowing the values of a patient’s parameters would allow prediction of treatment outcomes for any CD4:CD8 ratio using the model. In practice, most of these parameters, although biologically meaningful, are impractical or impossible to measure before treatment. However, some of the most sensitive parameters ($N$, $k_{b2}$, $k_{e}^{CD8}$, $k_{b1}$, $B50_{CD8}$) may be approximated as follows. (Rather than proposing a complete clinical workflow, this subsection demonstrates, using well-established experimental assays and clinically accessible measurements, how a subset of the most influential parameters identified by sensitivity analysis can be approximated in practice and mapped onto the model. The goal is to provide a concrete proof-of-concept for measurement-informed modeling, not a finalized protocol for clinical deployment.)
@@ 3.4 Neural Networks for Treatment Response Prediction
If the five aforementioned parameters can be measured, they can be used as inputs to the model, while all other parameters are set to typical patient values, allowing simulation of treatment outcomes.
To assess the robustness of model-based response classification to parameter uncertainty, we introduced additive Gaussian noise to the subset of five key patient-specific parameters identified by sensitivity analysis ($N$, $k_{b1}$, $k_{b2}$, $k_{e}^{CD8}$, and $B50_{CD8}$). For each virtual patient, noisy parameter values were generated by sampling from normal distributions centered at the true parameter values, with standard deviations equal to 5%, 10%, 25%, or 50% of the biologically feasible parameter ranges reported by Kirouac et al. (2023). These noise levels are not intended to model any specific experimental measurement process. Rather, they are designed to probe the limits of predictive performance under increasing uncertainty, including regimes in which parameter imprecision is large relative to biological variability (e.g. due to sparse, indirect, or noisy clinical measurements). This analysis therefore evaluates how sensitive mechanistic predictions are to parameter accuracy, and identifies conditions under which direct model-based classification may break down. Each patient was simulated under multiple treatment scenarios (no treatment, and CAR-T cell therapies with varying CD4:CD8 ratios) using both the true and noisy parameter sets. We classified treatment outcomes (complete response vs. non-response) for both parameter sets and assessed the consistency between them to quantify the model’s robustness.
Noise (%) Sensitivity (%) Specificity (%) Precision (%) Accuracy (%)
5 84.1 62.4 69.1 73.3
10 78.1 60.9 66.6 69.5
25 77.3 58.2 64.9 67.8
50 75.7 51.5 55.4 62.3
Table 3: Robustness of treatment response classification to parameter uncertainty. Average sensitivity (TPR), specificity (TNR), precision, and accuracy are reported across all tested conditions (No treatment; CD4 fraction = 0.0, 0.5, 0.8, and 1.0) as a function of increasing Gaussian noise applied to selected model parameters (5–50% of parameter range). Results correspond to simulations of $n=5000$ virtual patients.
As parameter uncertainty increases, the mechanistic model’s ability to correctly classify treatment response (CR vs. NR) deteriorates rapidly (Table 3). This degradation would be most consequential in regions of parameter space near any CR/NR decision boundary, where small errors can flip predicted outcomes. Notably, these are precisely the regimes in which predictive confidence is most clinically relevant: patients whose outcomes are sensitive to small changes in immune expansion or antigen dynamics are those for whom treatment selection and stratification are most challenging. In the remainder of this section, we therefore focus on the noisiest setting (50% of the biological parameter range) as a worst-case scenario to assess predictive limits under high uncertainty.
Treatment Sensitivity (%) Specificity (%) Precision (%) Accuracy (%)
NoTx 73.4 54.7 53.7 62.5
CD4frac 0.0 77.0 49.6 56.5 62.2
CD4frac 0.5 77.4 49.5 56.7 62.3
CD4frac 0.8 76.1 50.6 55.5 62.1
CD4frac 1.0 74.6 53.0 54.5 62.3
@@
Table 4: Classification performance of the tumor-immune ODE model under parameter noise ($n=5000$) across CD4:CD8 treatment conditions, with Gaussian noise applied at 50% of the biological parameter ranges reported by Kirouac et al. (2023). Sensitivity (true positive rate), specificity (true negative rate), precision (positive predictive value), and accuracy are reported.
@@
Metrics such as true positive rate (sensitivity) and true negative rate (specificity) were computed across all treatment conditions to summarize performance under this noisiest condition (Table 4). The poor performance of the model under noise, particularly specificity and precision, indicate that parameter measurement noise would make it difficult to reliably predict treatment outcomes. To address this challenge, we implemented a feed-forward neural network to train the model on this noisy data. The goal was to leverage the network’s ability to learn patterns from noisy inputs and improve predictive ability, thus mitigating the impact of measurement noise on the treatment outcome predictions.
@@
We trained a multi-label feed-forward neural network to predict, for each virtual patient, whether the mechanistic model yields a complete response (CR) or non-response (NR) under five treatment conditions simultaneously (one output per treatment). Inputs were the patient’s noisy values for five model parameters; targets were the corresponding noiseless simulation outcomes (CR vs. NR). Each input variable was standardized by subtracting the training-set mean and dividing by the training-set standard deviation to place features on a comparable scale. Virtual patient parameters in this study are sampled independently across dimensions, so the neural network cannot exploit cohort-specific parameter correlations to implicitly denoise inputs. Consequently, improvements in predictive performance reflect recovery of genuine signal under Gaussian noise rather than overfitting to spurious correlation structure. (Identifying and incorporating realistic biological correlations between parameters remains an important direction for future empirical work.)
@@
Data were split into a held-out test set (20%) and, from the remaining 80%, a validation set (20% of the remainder; 16% overall) and a training set. Training ran for up to 200 epochs with early stopping based on validation balanced accuracy at a fixed 0.5 threshold, restoring the best checkpoint. After training, decision thresholds were calibrated per treatment by sweeping a grid $t\in[0.01,0.99]$ on the validation set and selecting the threshold that maximized validation balanced accuracy. These validation-optimized thresholds were then applied once to the test set, and we report per-treatment sensitivity, specificity, precision, and accuracy (Table 5).
Treatment Sensitivity (%) Specificity (%) Precision (%) Accuracy (%)
NoTx 66.0 66.5 66.3 66.3
CD4frac 0.0 73.5 60.8 65.2 67.2
CD4frac 0.5 72.5 60.7 64.9 66.6
CD4frac 0.8 74.2 57.4 63.5 65.8
CD4frac 1.0 77.3 54.8 63.1 66.0
Table 5: Classification performance of the neural network on the test set (20 % of $n=5000$) across CD4:CD8 treatment conditions, with Gaussian noise applied at 50% of the biological parameter ranges reported by Kirouac et al. (2023). Sensitivity (true positive rate), specificity (true negative rate), precision (positive predictive value), and accuracy are reported.
Model Sensitivity (%) Specificity (%) Precision (%) Accuracy (%)
Neural network 72.7 60.0 64.6 66.4
Naïve baseline 64.8 35.2 64.9 50.1
Table 6: Average classification performance of the neural network compared with a naïve CR-rate baseline, averaged across the tested conditions (No treatment; CD4 fraction = 0.0, 0.5, 0.8, and 1.0) and trained on data with substantial Gaussian noise (standard deviation of 50% of the biological parameter range). Average sensitivity (true positive rate), specificity (true negative rate), precision (positive predictive value), and accuracy are reported. Results correspond to performance on the test set (20% of $n=5000$ virtual patients).
@@
As epochs increase, the training loss steadily decreases, while the validation loss drops at first and then plateaus; stopping training near the point where the validation loss plateaus (Fig. 7) helps prevent overfitting; i.e., it keeps the model from learning dataset-specific noise and preserves generalization to unseen data. In Fig. 8, the validation balanced accuracy rises quickly over the early epochs and then levels off, indicating that the network is learning useful signal and then converging. The plateau (with only small fluctuations) suggests diminishing returns from additional training rather than continued improvement, consistent with a stable model that generalizes to held-out data. Because balanced accuracy averages sensitivity and specificity, the upward trend reflects gains in both detecting responders and correctly rejecting non-responders despite class imbalance. Consistent with this, comparison to a naïve CR-rate baseline (Table 6) shows that the network’s performance gains are not driven by learning a global response bias, but by improved discrimination, most notably higher specificity and overall accuracy. Because the naïve classifier predicts CR at a fixed rate matched to the empirical prevalence, its precision is necessarily close to its sensitivity and reflects population frequency rather than true discriminative power. The superior performance of the neural network, therefore, indicates recovery of predictive signal beyond population-level statistics.
Although the neural network exhibits slightly reduced sensitivity relative to direct model-based classification in some treatment conditions, it substantially improves specificity, precision, and overall accuracy under high parameter uncertainty. This reflects a more conservative but more reliable classification strategy, reducing false-positive predictions while maintaining comparable true-positive detection. Consequently, the neural network achieves markedly higher balanced accuracy, indicating superior performance in the presence of severe parameter noise. Importantly, this shift toward higher precision and specificity may be clinically valuable, since false-positive responder predictions can lead to overly optimistic treatment selection and misguided stratification in the settings where uncertainty is highest.
We also used SHAP analysis to interpret feature importance across treatments (Figure 9(a)). Consistent with prior sensitivity analyses, variability in $N$ was the most influential factor overall (Figure 9(b)). The relative importance of the remaining parameters ($k_{b2}$, $k_{e}^{CD8}$, $k_{b1}$, $B50_{CD8}$) varied slightly across local, PRCC, and Sobol analyses, which is reflected in the SHAP ranking not exactly matching the Sobol order (Figure 9(c)). This indicates that while $N$ largely drives system behavior, interactions among the other parameters likely contribute nonlinearly to treatment outcomes. These results illustrate the value of combining mechanistic modeling with machine learning to better understand complex parameter effects under uncertainty.
@@ 4 Discussion
In this study, we extended the CAR-T cell therapy model proposed by Kirouac et al., which models antigen-regulated transitions between memory, effector, and exhausted T cell states, to explicitly model CD4+ and CD8+ T cell subtypes and their distinct roles in tumor-immune dynamics. As far as we know, our model is the first to differentiate between CD4+ and CD8+ CAR-T cells, allowing investigation of the synergistic interactions between these subtypes and their impact on treatment outcomes. Importantly, we do not view our extended model as a replacement for the original Kirouac et al. framework, but rather as a mechanistically interpretable extension that preserves its structure while expanding its biological scope.
@@
We use the extended model to investigate the effects of varying CD4:CD8 ratios on CAR-T cell efficacy and tumor control. In particular, the 1:1 CD4:CD8 ratio was examined because defined-composition CAR-T products with approximately equal CD4+ and CD8+ fractions have been investigated clinically and shown to yield improved efficacy relative to CD8-only formulations (Turtle et al., 2016). Our results demonstrate that a balanced 1:1 ratio of CD4+ to CD8+ CAR-T cells leads to more effective tumor elimination and CAR-T cell expansion on average (compared with a nonspecific CD4:CD8 ratio) but not always. As such, an important next step would be identification of the patient factors that determine whether a 1:1 CD4:CD8 treatment ratio or an unspecified ratio yields better therapeutic outcomes. Machine learning can be employed to analyze virtual patient simulation results to identify these factors. Identification of patient subgroups that are most likely to benefit from a specific CD4:CD8 ratio can enhance patient stratification and allow more personalized treatment. However, predictability of treatment response is dependent on limitations in quantifying patient-specific differences.
This highlights the distinction between mechanistic insight and predictive utility. Although the extended Kirouac et al. framework exhibits nonlinear outcome transitions under specific parameter variations, these behaviors depend strongly on patient-specific parameter combinations and are not universal; therefore, we used sensitivity analyses to systematically identify the parameters most influential for clinically relevant outcomes. We identified five such key parameters, and explored the effects of noise in patient-specific measurements of these parameters. Importantly, the noise levels considered here are intended as stress tests of predictive reliability rather than realistic models of experimental measurement error. While these noise levels may exceed typical experimental variability, substantial uncertainty can arise when parameters are inferred indirectly or estimated from sparse clinical data.
@@
We demonstrate that although noise in parameter measurements hinders direct predictions of treatment outcomes using the mathematical model, machine learning provides a potential solution for identifying possible treatment responders. We trained a neural network on noisy parameter sets and corresponding simulation outcomes to predict virtual patient responses, achieving improved overall performance for direct predictions from the mathematical model. It is important to note, however, that this recovery is partial and does not eliminate the need for accurate parameter measurement; rather, it highlights regimes where data-driven methods can complement mechanistic models when uncertainty is unavoidable. Additionally, the agreement between SHAP and mechanistic sensitivity analyses strengthens confidence in the model’s predictions and illustrates the complementarity of mechanistic and machine learning approaches.